OpenStudy (anonymous):

Based on the graph below, what is the total number of solutions to the equation f(x) = g(x)? graph of function f of x equals negative 2 multiplied by x plus 3 and graph of function g of x equals x cubed plus 4 multiplied by x squared minus x minus 4 One Two Three Four

OpenStudy (anonymous):

please helppp

OpenStudy (anonymous):

??

OpenStudy (anonymous):

ill give a medal

OpenStudy (anonymous):

@Adjax

OpenStudy (anonymous):

\(f(x)=2x+3\) or \(f(x)=2(x+3) ?\) \(g(x)=x^3+4x^2-x-4\) Can you verify the equations for me?

OpenStudy (anonymous):

What equations? how

OpenStudy (anonymous):

I'm really bad at math lol

OpenStudy (anonymous):

Like for f(x) I am not sure which one is correct ... Look it cld be either equation

OpenStudy (anonymous):

hold on ill attach the exact graph for you

OpenStudy (anonymous):

ya that would be great ... cuz what you wrote above is slightly ambiguous

OpenStudy (anonymous):

haha yea, give me a min

OpenStudy (anonymous):

wait actually can you help with a different question i already got the answer

OpenStudy (anonymous):

ya sure -.-

OpenStudy (anonymous):

Im assuming this is some online course .. k go for it

OpenStudy (anonymous):

yea flvs lol

OpenStudy (anonymous):

For the graph below, what should the domain be so that the function is at least 600? graph of y equals minus 2 times the square of x plus 40 times x plus 600 0 ≤ x ≤ 20 -10 ≤ x ≤ 30 x ≥ 0 all real numbers

OpenStudy (anonymous):

hey there are 2 complex and one real solution for that I guess

OpenStudy (anonymous):

so total of 3

OpenStudy (anonymous):

Its stupid flvs courses ...

OpenStudy (anonymous):

thank you :) that was the first question right?

OpenStudy (anonymous):

yeah..The pge just loaded at the moment

OpenStudy (anonymous):

Ok look at the graph ...

OpenStudy (anonymous):